OpenStudy (anonymous):

8x/5x+3=10x-2/7x find x

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@dtan5457 @campbell_st @confluxepic @sleepyjess

OpenStudy (dtan5457):

\[\frac{ 8x }{ 5x+3 }=\frac{ 10x-2 }{ 7x }\]

OpenStudy (dtan5457):

like this?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

Cross multiply

OpenStudy (dtan5457):

Equate them together and form a quadratic, then solve using quadratic formula

OpenStudy (anonymous):

does x=1/6 @dtan5457

OpenStudy (dtan5457):

No. It's gonna be in the form of the quadratic formula.

OpenStudy (dtan5457):

When you cross multiply, what do you get?

OpenStudy (anonymous):

(10x-2)(5x+3)=56x

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

8x times 7x=56x^2

OpenStudy (dtan5457):

Solve for x using the quadratic formula 56x^2=50x^2+20x-6 isolate 6x^2-20x+6 factor x and solve..

OpenStudy (anonymous):

is that a minus sign or is it suppose to be an equal sign? @dtan5457

OpenStudy (dtan5457):

Minus, I moved all terms to one side, it should technically be 6x^2-20x+6=0

OpenStudy (anonymous):

@dtan5457 i cant figure out what x is

OpenStudy (dtan5457):

Not sure how to factor it? You can try using the quadratic formula, if you want.

OpenStudy (dtan5457):

Or do you want to learn how to factor?

OpenStudy (anonymous):

whats the quadratic factor

OpenStudy (dtan5457):

Quadratic formula is \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] You can try this once, but if it fails, I'll show you my method.